OpenStudy (babystar16):

airline fares for a flight from dallas to austin are $30 for first class and $25 for tourist class. if a flight had 52 passengers who paid a total of $1360, how many first class passengers were on the trip?

zepdrix (zepdrix):

Let F be the number of number of first class passengers Let T be the number of tourist class passengers. $30 per (F)first class + $25 per (T)tourist class = $1360. \[30F + 25T = 1360\]Total number of passengers is 52. \[F + T = 52\] So we have a system of 2 equations, with 2 unknowns. To solve this, we can use either substitution or elimination. Are you familiar with either method? c:

OpenStudy (babystar16):

yes i just had problems solving it

zepdrix (zepdrix):

\[F+T=52\]\[T=52-F\] Plugging this into the other equation gives us. \[30F+25T=1360\]\[30F+25(52-F)=1360\]\[30F+1300-25F=1360\]\[5F=60\]\[F=12\] So that's how we would find F. Confused about any of the steps? :o

zepdrix (zepdrix):

This is the substitution method :3